OpenStudy (prince1342):

Simplify. 4 square root of 2 end root plus 7 square root of 2 end root minus 3 square root of 2

OpenStudy (prince1342):

Ill medal anyone who can eplain to me how to complete it

OpenStudy (anonymous):

\[4\sqrt{2}\sqrt{7}\sqrt{2}-3\sqrt{2}\] Not quite sure if this is what you're looking for based on how to set it up.